OpenStudy (anonymous):

simplify x^2-25/x^2-3-10

OpenStudy (zzr0ck3r):

should that say -3x-10 on the bottom?

OpenStudy (mertsj):

\[\frac{x^2-25}{x^2-13}\]

OpenStudy (mertsj):

That is simple form.

OpenStudy (anonymous):

yes sorry x^2-3x-10

OpenStudy (zzr0ck3r):

x^2-25 = (x-5)(x+5) x^2 - 3x - 10 = (x+5)(x-2)

OpenStudy (mertsj):

\[\frac{(x-5)(x+5)}{(x-5)(x+2)}\]

OpenStudy (zzr0ck3r):

does that help?

OpenStudy (mertsj):

Cancel the common factor.

OpenStudy (anonymous):

Thank you guys this helps

OpenStudy (anonymous):

factor cancel and tell your math teacher (gently) that there is no such mathematical operation as "simplify" what he or she meant is "reduce to lowest terms"

OpenStudy (mertsj):

yw

OpenStudy (zzr0ck3r):

lol